# Spectra of Magnetoroton and Chiral Graviton Modes of Fractional Chern Insulator

### v2

@@ Abstract
Figure 4: Chiral graviton mode (CGM) inside FCI phase ($V_{1}=4,V_{2}=1$) and its tunability. (a) and (b) are the dynamic quadrupolar density correlations $g^{+}$ ($S=2$) and $g^{-}$ ($S=-2$), respectively. The chiral $S=-2$ graviton extends around $\Gamma$ point (denoted by the red dashed ellipse). In (c) and (d), we revert the magnetic flux $\phi$ such that the Chern number for lower band is $C=-1$ and calculate $g^{+}$ and $g^{-}$ as in (a) and (b). The chirality of the graviton mode is clearly reversed.
@@
Fig. 4 (a) and (b) are the $g^{+}(\mathbf{k},\omega)$ and $g^{-}(\mathbf{k},\omega)$ at $V_{1}=0,V_{2}=0$ inside the FCI phase. One sees a strong charge neutral response close to the momentum $\Gamma$ at $(S=-2)$ sector. This is also the momenta where the CGM was proposed in the FQHE literature [35, 68] and experiment [43]. One more interesting analysis we did here is to switch the Chern number and chirality of the flat band in our model, this is similar with the observation of the CGM in the FQHE experiment that hole doping FQHE state ($\nu=2/3$) have opposite chirality compared with it’s electron ($\nu=1/3$) counterpart. As shown in Fig. 4 (c) and (d), once we revert the flux directions in the in the $t^{\prime}$-term in Eq. (1), at the same interaction strength of FCI phase, the graviton appears near the $\Gamma$ point in $g^{+}(\mathbf{k},\omega)$. The CGM at finite momentum could lead to spatial modulation of internal quantum metric and more easily captured by experiment. We think such tunability of the chirality of the graviton in the FCI system could be potentially observed in the circularly polarized resonant inelastic light scattering experiment [69, 43].
@@
Discussion.— In this work, we observe the clear spectral properties of the magnetoroton and chiral graviton, as the collective charge neutral excitations in FCI system. We discover that the roton minimum can go soft and give rise to a CDW phase (with the corresponding wavevector), while during the process, the spectral weight of the roton excitation become sharper and more concentrated but the single-particle gap remains large. The enhanced spectral signature and the potentially associated enhanced charge compressibility, are expected to be measured in STS and quantum capacitance experiments. Moreover, we not only show the graviton mode is chiral for a given FCI, but also directly visualize the different chiralities of the CGM of FCIs by tuning the sign of Hall conductance and such novel phenomena can potentially be seen via the resonant inelastic light scattering, as has recently been seen in FQHE experiment [43].

### v3

@@ Abstract
Figure 4: Chiral graviton mode (CGM) inside FCI phase and its tunability. (a) and is the dynamic quadrupolar density correlations $g^{-}$ ($S=-2$) measured at $V_{1}=4,V_{2}=1$. The chiral $S=-2$ graviton extends around $\Gamma$ point. In (b), we revert the magnetic flux $\phi$ such that the Chern number for the lower band is $C=-1$ and calculate $g^{+}$ as in (a). (c) and (d) are the same measurement as (a) and (b), but on $1/3$ FCI with $V_{1}=1.5,V_{2}=0$, where the chirality reversion is also observed
@@
Fig. 4 (a) is the $g^{-}(\mathbf{k},\omega)$ at $V_{1}=4,V_{2}=1$ inside the $1/2$ bosonic FCI phase. One sees a strong charge-neutral response close to the momentum $\Gamma$ at $(S=-2)$ sector with energy twice that of the magnetoroton. This is also the momentum where the CGM was proposed in the FQHE literature [17, 63, 64, 65, 38], and experiment [18]. One more interesting analysis we did here is to switch the Chern number and chirality of the flat band in our model, this is similar with the observation of the CGM in the FQHE experiment that hole doping FQHE state ($\nu=2/3$) have opposite chirality compared with it’s electron ($\nu=1/3$) counterpart or flip the direction of the magnetic field. As shown in Fig. 4 (b), once we revert the flux directions in the in the $t^{\prime}$-term in Eq. (1), at the same interaction strength of FCI phase, the graviton appears near the $\Gamma$ point in $g^{+}(\mathbf{k},\omega)$. The result for $1/3$ fermionic FCI with $V_{1}=1.5,V_{2}=0$ is shown in Fig. 4 (c)(d), the spectrum peak is located at the $\Gamma$ point, and the chiral mode switching is observed upon reversing the flux directions. However, the chiral signal is weaker and of distinct shape compared with the bosonic case. For bosonic FCI, we think such tunability of the chirality of the graviton in the FCI system could be potentially observed in the circularly polarized resonant inelastic light scattering experiment  [66, 18]. For fermionic FCI, we also expect the tunability of the chirality for CGM.
We also trace the evolution of the spectrum weight upon increasing cylinder length. The result is shown in Section II of the SM[42]. The spectrum peak gradually declines at larger system sizes. This trend has been observed on both FQH[12] and FCI states[65], which is attributed to the increase of density of states at larger systems, introducing more scattering channels and consequently lowering the lifetime of CGM[65]. Although our phenomenological operator has a much wider momentum spread compared with the bosonic case. We leave the clearer identification of the CGM in fermionic FCI models and its lifetime in the thermodynamic limit for future investigations.
@@
Discussion and conclusions.— In this work, we observe the clear spectral properties of the magnetoroton and chiral graviton, as the collective charge-neutral excitations in FCI systems. We discover that the roton minimum can go soft and give rise to a CDW phase (with the corresponding wavevector), while during the process, the spectral weight of the roton excitation becomes sharper and more concentrated, while the single-particle gap remains large. The enhanced spectral signature is expected to be measured in Raman scattering. Moreover, we not only show that the graviton mode is chiral for a given FCI state, but also directly visualize the different chiralities of the CGM of FCIs by tuning the sign of the Hall conductance and such novel phenomena can potentially be seen via the resonant inelastic light scattering, as has recently been seen in the FQH experiment [18].